Several intelligence tests follow a normal distribution with a mean of 100 and a standard deviation of 15.


a. Determine the percentage of the population that would obtain a score between 95 and 110.

b. For a population of 2,500, how many are expected to have a score above 125?


1
Expert's answer
2021-11-08T15:36:59-0500

a. "P(95<X<110)"

"P(95<X<110)=P(\frac{95-100}{15}<Z<\frac{110-100}{15})"

"=P(Z<0.6667)-P(Z< -0.3333)"

"=0.7475-0.3694=0.3781"

Thus, 37.81% of the population would obtain scores between 95 and 110


b. "P(X>125)=1-P(X<125)"

"=1-P(Z<\frac{125-100}{15})=1-P(Z<\frac{5}{3})=1-0.9522=0.0478"

For a population of 2,500,

"2500\times0.0478=119.5"

Thus, 120 people out of 2500 will score above 125.